Safawi kicked a ball with an initial speed of 20 ms-1 at 30º angle. If the ball experienced a constant vertical acceleration of -9.81 ms-2, calculate the ball’s maximum height and distance. Explain how soccer player could apply the force of summation theory to optimize his kicks performance. 


1
Expert's answer
2022-02-09T10:50:17-0500

"v_0=20\ \text{m/s}^2"

"\alpha =30"


The maximum height "H=\frac{v_0^2\sin^2\alpha}{2g}"

The distance "L=\frac{v_0^2\sin 2\alpha}{g}"


"H=\frac{20^2\cdot \sin^2 30^\circ }{2\cdot 9.81}=\frac{100}{19.62}=5.09\ \text{m}"


"L=\frac{20^2\cdot \sin 60^\circ }{9.81}=\frac{200\sqrt{3}}{9.81}=35.31\ \text{m}"


Answer: maximum height = 5.09 m, distance = 35.31 m. A soccer player can apply the force of summation theory to maximize the ball’s distance. It depends on an initial speed an angle.